# Electronic and magnetic properties of graphene-fluorographene nanoribbons: controllable semiconductor-metal transition

### v1

@@ VI Mean field Anderson-Hubbard Model from Wannier calculations
Figure 19 compares the paramagnetic (PM) and ferromagnetic states of the $40$-$\alpha\beta$ system. In the PM state, panel (a), electronic states near $\mathrm{k}=\frac{\pi}{a}$ are fully occupied. In contrast, the FM state, panel (b), exhibits spin-splitting, with one spin band filled and the other empty near $\mathrm{k}=\frac{\pi}{a}$. This difference reflects the presence of spin polarization along the $\alpha$ edge in the latter. Finally, Fig. 19(c) shows a near-constant energy difference ($\delta E$) between FM and PM states for all system sizes. This is due to the localized nature of the edge-state interactions responsible for the magnetic ordering.
@@ VII Summary
In this work we studied the electronic and magnetic properties of graphene channels embedded within fluorographene, focusing on two distinct interfaces: the fully fluorinated zigzag chain ($\alpha$) and the half-fluorinated zigzag chain ($\beta$).
In the case of the $n$-$\alpha\alpha$ channels we found that they exhibit similarities to pristine zigzag graphene nanoribbons: they display antiferromagnetic ordering, semiconducting behavior, and a decreasing energy gap with increasing channel width. The $n$-$\alpha\beta$ systems, on the other hand, are ferromagnetic and undergo a semiconductor-to-metal transition with increasing channel width. In the semiconducting phase, spin polarization is present at both edges. For wider channels, dominant spin polarization persists only at the $\alpha$ edge, with vanishing polarization at the $\beta$ edge.
Our results demonstrate that selective fluorination of graphene edges offers an opportunity for tunable electronic and magnetic properties of significant interest for spintronic applications. This strategy provides a pathway for engineering semiconductor properties without altering the graphene channel width itself. For example, inducing metallicity in graphene nanoribbons has been a subject of interest recently [9], our results show that by tuning the interface and by fabricating $\alpha\beta$ nanoribbons, one can also obtain metallic nanoribbons. Graphene nanoribbons are promising devices with potential uses in microelectronics, spintronics, quantum computing and topological electronics. Our findings highlight the potential of selective fluorination for tailoring graphene’s properties, paving the way for such developments.
@@ VIII Acknowledgments
We acknowledge financial support from the ANPCyT-FONCyT (Argentina) under grants PICT 2018-1509 and PICT 2019-0371, from CONICET grant PIP 2022-2024 (11220210100041CO), and SeCyT-UNCuyo grant 06/C053-T1. GU thanks Nathan Goldman for his hospitality at the Université Libre de Bruxelles (ULB). *
@@ Appendix A Parameters used in the Anderson-Hubbard model
We describe here the procedure we use to obtain the different parameters of the mean field Anderson-Hubbard model described in the main text and schematically depicted in Fig. 20. Figs. 20(a) and 20(b) highlight the locations of the ${\pi}$ (black and red dots) and the (${\mathcal{A}}$) (gray dots) orbitals, used for describing the graphene and fluorographene regions, respectively. Fig. 20(c) shows the underlying hexagonal lattice containing one Wannier orbital per vertex. The dashed lines defines the unit cell. Periodic boundary conditions and translation invariance are assumed along the $y$ and $x$ axes, respectively.
Figure 20: (Color online) Crystal structure and Anderson-Hubbard model for the $6$-$\alpha\alpha$ and the $6$-$\alpha\beta$ nanoribbons. (a) and (b) illustrate the key structural difference: the number of graphene sites (black and red circles) within the unit cell. The $6$-$\alpha\alpha$ system has $12$ sites, whereas the $6$-$\alpha\beta$ system has $13$. This variation significantly alters the band structures, as discussed previously. Gray circles correspond with the fluorographene sites. (c) Depicts a simplified lattice representation of the Hamiltonian (Eq. 2). The hexagonal lattice shows the unit cell (enclosed by dashed lines) and numbered atomic sites. The model considers one orbital per lattice vertex and periodic boundary conditions in both directions. Additionally, it includes hopping terms (notated as $t_{n}$ with $n=1,2,3,4,5,6$) up to the sixth-nearest neighbors.
Notice that, although both $\alpha\alpha$ and $\alpha\beta$ systems have $24$ sites per unit cell, the distribution of ${\pi}$ and ${\mathcal{A}}$ orbitals at the graphene-fluorographene interface differs. The graphene region contains $12$ sites in the $\alpha\alpha$ case and $13$ sites in the $\alpha\beta$ case.
The Hubbard interaction strength $U_{j}$ is self-consistently calculated from the local energies, eigenvalues, and eigenvectors (Eq. (4)) of the spin-resolved Wannier Hamiltonians ($\mathcal{H}_{W}^{\sigma}$). Figure 21 shows its spatial variation along the graphene channels. To simplify our model, we adopt a representative value of $U_{g}=3.45$ eV for all graphene sites as this is roughly the value obtained near the highly spin-polarized edges—except a slight deviation at the $\beta$ edge in the smallest ($4$-$\alpha\beta$) system.
@@
Figure 22: Local energy ($\varepsilon_{j}$) of the system calculated using Eq. (4). Upper panels correspond to the graphene channels and the lower panels show the behavior along the fluorographene region. Grey circles represent $\varepsilon_{j}$ values, which exhibits similar behavior near $\alpha$ and $\beta$ edges across different nanoribbon sizes. This allows us to reliably use these values to model wider systems inaccessible to DFT calculations. For larger systems, $\varepsilon_{j}$ within the central region can be described by interpolating the straight lines and employ the values at the edges (black open circles).
The site energies $\varepsilon_{j}$, obtained by fitting Eq. (5), are shown in Fig. 22. In the graphene channels (upper panels), the interface effect extends approximately $3$ sites away from $\alpha$-edges and up to $6$ sites from $\beta$-edges. The potential decreases towards the channel center. Conversely, within fluorographene regions (lower panels), the potential increases towards the center, with the interface effect extending roughly $3$ sites from both $\alpha$ and $\beta$ edges. Importantly, Fig. 22 reveals that the behavior of $\varepsilon_{j}$ near edges is consistent across different nanoribbon sizes. This allows us to model wider systems (intractable for DFT) by employing these values. For the central regions of larger systems, we interpolate $\varepsilon_{j}$ based on the linear trends observed in our smaller systems.
Our next step is to calculate the hopping matrix elements. In Fig. 15, we show the importance of including hopping terms up to second-nearest neighbors for $\alpha\alpha$ interfaces and fifth-nearest neighbors for $\alpha\beta$ interfaces to achieve DFT-level accuracy near the Fermi level. This means that in $\alpha\alpha$ systems, hopping matrix elements will link orbitals in adjacent unit cells, while in $\alpha\beta$ systems they might connect orbitals across second-nearest neighbor unit cells. To calculate this hopping matrix elements we used the Wannier Hamiltonians for the $6$-$\alpha\alpha$ the $8$-$\alpha\beta$ nanoribbons.
@@
where $\zeta_{l}$ represents the $\pi$ or $\mathcal{A}$ orbital depending on the site. The spin averaging eliminates a weak spin-dependence of matrix elements. Let us first consider the matrix elements for sites within the graphene ($t$) and fluorographene ($\tau$) channels, away from interfaces. Since the system is slightly anisotropic, with the goal of reducing the number of parameters, we define the average $m$-th order matrix element at each site, say $l$, as
@@
where $Z_{ml}$ is the number of $m$-th nearest neighbors of site $l$ and the sum on $i$ runs on those neighbors.
@@
Table 3: Average $n$-th order matrix elements at the center of graphene ($\langle^{n}t_{l}\rangle$) and fluorographene ($\langle^{n}\tau_{l}\rangle$ ) channels, calculated using Eq. (7).
As expected, $\langle^{1}t_{l}\rangle$ and $\langle^{1}\tau_{l}\rangle$ have the largest magnitudes, as they encode the essential information about the crystal structure. In particular, their values are significantly altered by the presence of interfaces as discussed below.
On the other hand, as mentioned in the main text, a successful model requires including matrix elements beyond first-nearest neighbors to accurately capture interface effects. Figure 23 depicts the relevant matrix elements near the $\alpha$ and $\beta$ interfaces. Black dots represent graphene $\pi$ orbitals, while gray dots represent fluorographene $\mathcal{A}$ orbitals. Panels 23(a) and 23(c) show first- and second-nearest neighbor interactions for $\alpha$ and $\beta$ interfaces, respectively, Panels 23(b) and 23(d) display third- through fifth-nearest neighbor interactions, while
@@
Table 5: Matrix elements corresponding to the hoppings near the $\alpha$ interface depicted in Figs. 23(a) and 23(b). These elements include 1st through 5th nearest neighbors and were extracted from Wannier analysis of the 8-$\alpha\beta$ system.
@@
Figure 23: (Color online) Matrix elements between interface-adjacent orbitals in graphene (black dots) and fluorographene (orange dots). Panels (a) and (b) depict the $\alpha$ interface; panels (c) and (d) depict the $\beta$ interface. (a) and (c): first- and second-nearest neighbor interactions. (b) and (d): third- through fifth-nearest neighbor interactions.
@@
The spatial dependence of the first-nearest neighbor matrix elements, $t_{j+1,j}$, along the graphene channel is also relevant. Figures 24(a) and 24(b) show $t_{j+1,j}$ vs. orbital position ($j$) for the $6$-$\alpha\alpha$ and $8$-$\alpha\beta$ systems, respectively, as obtained from Eq. (6). A distinct separation emerges between even and odd $j$ values within the central region. Matrix elements for even $j$ exhibit a larger magnitude (more negative) than those for odd $j$, see Fig. 20(c) for orbital labeling. This behavior can be understood by the presence of internal stress within the graphene-fluorographene superlattice: fluorographene’s larger lattice parameter stretches the graphene along the longitudinal axis ($\hat{x}$-direction). In fact, Fig. 20(c) shows that when $j$ is even the hopping $t_{j+1,j}$ involves orbitals separated along the $\hat{y}$ direction while when $j$ is odd they are separated along the $\frac{\sqrt{3}}{2}\hat{x}\pm\frac{1}{2}\hat{y}$ direction. In both cases, the distance between orbitals should be the same for a system without internal stress. However, in the superlattice of alternating graphene-flurographene nanoribbons, the graphene channel is elongated along the longitudinal $\hat{x}$ direction, which increases the distance between orbitals that were initially in the $\frac{\sqrt{3}}{2}\hat{x}\pm\frac{1}{2}\hat{y}$ direction (odd $j$) if we compare it with the separation of orbitals along the $\hat{y}$ direction (even $j$), leading to weaker hopping terms.
@@
Figure 24: (Color online) Variation of first-nearest neighbor matrix elements $t_{(j+1,j)}$ as a function of orbital position $j$. (a) 6-$\alpha\alpha$ system and (b) 8-$\alpha\beta$ system. (c) $t_{(j+1,j)}$ as a function of the inter-orbital distance ($d_{(j+1,j)}$). Lines represent the fitted functional dependence (indicated in (c)), with data points corresponding to the different system sizes.
@@
To quantify this effect, Fig. 24(c) plots $t_{j+1,j}$ against the inter-orbital distance ($d_{j+1,j}$) obtained from relaxed DFT geometries. The exponential dependence is apparent from the figure [30]. We fit this dependence using $t(d)=t_{0}\exp{(\gamma_{l}(\frac{d}{a_{0}}-1)})$, obtaining different parameters for even and odd $j$, reflecting the stress-induced anisotropy: for $j$ odd (black symbols) $t_{0}=-2.88$ eV and $\gamma=-3.16$; for $j$ even (gray symbols) $t_{0}=-2.82$ eV and $\gamma=-2.17$. In both cases, $a_{0}=1.42$Å. Importantly, the value of $t_{j+1,j}$ near the edges deviate from the fitted trends of the central region. This is due to the local distortion induced by the interface.
@@
In summary, to model $n$-$\alpha\alpha$ systems, we use the following system-size independent scheme for the first-nearest neighbor hopping terms
@@
The corresponding parameter for $n$-$\alpha\beta$ systems are
@@
In this case, the hopping term $t_{j+1,j}$ in the bulk of the channel has a constant value for $j$ odd, ${}^{1}t_{c}=-2.753$ eV, and a linear dependence ${}^{1}\mathrm{t}_{t}(j)={{}^{1}}\mathrm{t}_{\alpha c}+\frac{{}^{1}\mathrm{t}_{\beta c}-^{1}\mathrm{t}_{\alpha c}}{2n-8}(j-4)$ for even $j$, with ${}^{1}t_{\beta c}=-2.854$ eV and ${}^{1}t_{\alpha c}=-2.804$ eV, in accordance with Fig. 24(b).

### v2

@@ VI Mean field Anderson-Hubbard Model from Wannier calculations
Figure 19 compares the paramagnetic (PM) and ferromagnetic states of the $40$-$\alpha\beta$ system. In the PM state, [Fig. 19(a)], electronic states near $\mathrm{k}=\frac{\pi}{a}$ are fully occupied. In contrast, the FM state, [Fig. 19 (b)], exhibits spin splitting, with one spin band filled and the other empty near $\mathrm{k}=\frac{\pi}{a}$. This difference reflects the presence of spin polarization along the $\alpha$ edge in the latter. Finally, Fig. 19(c) reveals a sharp increase in $\delta E$ for small system sizes, attributed to the semiconductor-metal transition depicted in Fig.  6. For larger systems, the energy difference ($\delta E$) between the FM and PM states remains nearly constant. This is due to the localized nature of the edge-state interactions responsible for the magnetic ordering.
@@ VII Summary
In this work, we studied the electronic and magnetic properties of graphene channels embedded within fluorographene, focusing on two distinct interfaces: the fully fluorinated zigzag chain ($\alpha$) and the half-fluorinated zigzag chain ($\beta$).
In the case of the $n$-$\alpha\alpha$ channels, we found that they exhibit similarities to pristine zigzag graphene nanoribbons: they display antiferromagnetic ordering, semiconducting behavior, and a decreasing energy gap with increasing channel width. The $n$-$\alpha\beta$ systems, on the other hand, are ferromagnetic and undergo a semiconductor-to-metal transition with increasing channel width. In the semiconducting phase, spin polarization is present at both edges. For wider channels, dominant spin polarization persists only at the $\alpha$ edge, with vanishing polarization at the $\beta$ edge.
Our results demonstrate that selective fluorination of graphene edges offers an opportunity for tunable electronic and magnetic properties of significant interest for spintronic applications. This strategy provides a pathway for engineering semiconductor properties without altering the graphene channel width itself. For example, inducing metallicity in graphene nanoribbons has been a subject of interest recently [9]; our results show that by tuning the interface and by fabricating $\alpha\beta$ nanoribbons, one can also obtain metallic nanoribbons. Graphene nanoribbons are promising devices with potential uses in microelectronics, spintronics, quantum computing, and topological electronics. Our findings highlight the potential of selective fluorination for tailoring graphene’s properties, paving the way for such developments.
@@ VIII Acknowledgments
We acknowledge financial support from the ANPCyT-FONCyT (Argentina) under Grants No. PICT 2018-1509 and No. PICT 2019-0371, from CONICET Grant No. PIP 2022-2024 (11220210100041CO), and SeCyT-UNCuyo grant 06/C053-T1. G.U. thanks Nathan Goldman for his hospitality at the Université Libre de Bruxelles (ULB). *
@@ Appendix A Parameters used in the Anderson-Hubbard model
We describe here the procedure we use to obtain the different parameters of the mean field Anderson-Hubbard model described in the main text and schematically depicted in Fig. 20. Figures. 20(a) and 20(b) highlight the locations of the ${\pi}$ (black and red dots) and the ${\mathcal{A}}$ (gray dots) orbitals, used for describing the graphene and fluorographene regions, respectively. Figure 20(c) shows the underlying hexagonal lattice containing one Wannier orbital per vertex. The dashed lines defines the unit cell. Periodic boundary conditions and translation invariance are assumed along the $y$ and $x$ axes, respectively.
Figure 20: Crystal structure and Anderson-Hubbard model for the $6$-$\alpha\alpha$ and the $6$-$\alpha\beta$ nanoribbons. (a), (b) The key structural difference: The number of graphene sites (black and red circles) within the unit cell. The $6$-$\alpha\alpha$ system has $12$ sites, whereas the $6$-$\alpha\beta$ system has $13$. This variation significantly alters the band structures, as discussed previously. Gray circles correspond with the fluorographene sites. (c) A simplified lattice representation of the Hamiltonian [Eq. 2]. The hexagonal lattice shows the unit cell (enclosed by dashed lines) and numbered atomic sites. The model considers one orbital per lattice vertex and periodic boundary conditions in both directions. Additionally, it includes hopping terms (notated as $t_{n}$ with $n=1,2,3,4,5,6$) up to the sixth-nearest neighbors.
Notice that although both $\alpha\alpha$ and $\alpha\beta$ systems have $24$ sites per unit cell, the distribution of ${\pi}$ and ${\mathcal{A}}$ orbitals at the graphene-fluorographene interface differs. The graphene region contains $12$ sites in the $\alpha\alpha$ case and $13$ sites in the $\alpha\beta$ case.
The Hubbard interaction strength $U_{j}$ is self-consistently calculated from the local energies, eigenvalues, and eigenvectors [Eq. (4)] of the spin-resolved Wannier Hamiltonians ($\mathcal{H}_{W}^{\sigma}$). Figure 21 shows its spatial variation along the graphene channels. To simplify our model, we adopt a representative value of $U_{g}=3.45$ eV for all graphene sites as this is roughly the value obtained near the highly spin-polarized edges—except a slight deviation at the $\beta$ edge in the smallest ($4$-$\alpha\beta$) system.
@@
Figure 22: Local energy ($\varepsilon_{j}$) of the system calculated using Eq. (4). (a)-(d) The graphene channels and (e)-(h) the behavior along the fluorographene region. Gray circles represent $\varepsilon_{j}$ values, which exhibit similar behavior near $\alpha$ and $\beta$ edges across different nanoribbon sizes. This allows us to reliably use these values to model wider systems that are inaccessible to DFT calculations. For larger systems, $\varepsilon_{j}$ within the central region can be described by interpolating the straight lines and employ the values at the edges (black open circles).
The site energies $\varepsilon_{j}$, obtained by fitting Eq. (5), are shown in Fig. 22. In the graphene channels (upper panels), the interface effect extends approximately three sites away from $\alpha$ edges and up to six sites from $\beta$ edges. The potential decreases towards the channel center. Conversely, within fluorographene regions (lower panels of Fig. 22), the potential increases towards the center, with the interface effect extending roughly three sites from both $\alpha$ and $\beta$ edges. Importantly, Fig. 22 reveals that the behavior of $\varepsilon_{j}$ near edges is consistent across different nanoribbon sizes. This allows us to model wider systems (intractable for DFT) by employing these values. For the central regions of larger systems, we interpolate $\varepsilon_{j}$ based on the linear trends observed in our smaller systems.
Our next step is to calculate the hopping matrix elements. In Fig. 15, we show the importance of including hopping terms up to second-nearest neighbors for $\alpha\alpha$ interfaces and fifth-nearest neighbors for $\alpha\beta$ interfaces to achieve DFT-level accuracy near the Fermi level. This means that in $\alpha\alpha$ systems, hopping matrix elements will link orbitals in adjacent unit cells, while in $\alpha\beta$ systems they might connect orbitals across second-nearest neighbor unit cells. To calculate these hopping matrix elements we used the Wannier Hamiltonians for the $6$-$\alpha\alpha$ the $8$-$\alpha\beta$ nanoribbons.
@@
where $\zeta_{l}$ represents the $\pi$ or $\mathcal{A}$ orbital depending on the site. The spin averaging eliminates a weak spin dependence of matrix elements. Let us first consider the matrix elements for sites within the graphene ($t$) and fluorographene ($\tau$) channels, away from interfaces. Since the system is slightly anisotropic, with the goal of reducing the number of parameters, we define the average $m$th-order matrix element at each site, say $l$, as
@@
where $Z_{ml}$ is the number of $m$th-nearest neighbors of site $l$ and the sum on $i$ runs on those neighbors.
@@
Table 3: Average $n$th-order matrix elements at the center of graphene (GC, $\langle^{n}t_{l}\rangle$) and fluorographene (FG, $\langle^{n}\tau_{l}\rangle$ ) channels, calculated using Eq. (7).
As expected, $\langle^{1}t_{l}\rangle$ and $\langle^{1}\tau_{l}\rangle$ have the largest magnitudes, as they encode the essential information about the crystal structure. In particular, their values are significantly altered by the presence of interfaces, as discussed below.
On the other hand, as mentioned in the main text, a successful model requires inclusion of the matrix elements beyond first-nearest neighbors to accurately capture interface effects. Figure 23 depicts the relevant matrix elements near the $\alpha$ and $\beta$ interfaces. Black dots represent graphene $\pi$ orbitals, while gray dots represent fluorographene $\mathcal{A}$ orbitals. Figures 23(a) and 23(c) show first- and second-nearest-neighbor interactions for $\alpha$ and $\beta$ interfaces, respectively, Panels 23(b) and 23(d) display third- through fifth-nearest neighbor interactions, while
@@
Table 5: Matrix elements corresponding to the hoppings near the $\alpha$ interface depicted in Figs. 23(a) and 23(b). These elements include first- through fifth-nearest neighbors and were extracted from Wannier analysis of the 8-$\alpha\beta$ system.
@@
Figure 23: Matrix elements between interface-adjacent orbitals in graphene (black dots) and fluorographene (orange dots). (a), (b) The $\alpha$ interface; (c), (d) The $\beta$ interface. (a), (c) First- and second-nearest-neighbor interactions. (b), (d): Third- through fifth-nearest-neighbor interactions.
@@
The spatial dependence of the first-nearest neighbor matrix elements, $t_{j+1,j}$, along the graphene channel is also relevant. Figures 24(a) and 24(b) show $t_{j+1,j}$ vs orbital position ($j$) for the $6$-$\alpha\alpha$ and $8$-$\alpha\beta$ systems, respectively, as obtained from Eq. (6). A distinct separation emerges between even and odd $j$ values within the central region. Matrix elements for even $j$ exhibit a larger magnitude (more negative) than those for odd $j$; see Fig. 20(c) for orbital labeling. This behavior can be understood by the presence of internal stress within the graphene-fluorographene superlattice: fluorographene’s larger lattice parameter stretches the graphene along the longitudinal axis ($\hat{x}$ direction). In fact, Fig. 20(c) shows that when $j$ is even, the hopping $t_{j+1,j}$ involves orbitals separated along the $\hat{y}$ direction, while when $j$ is odd, they are separated along the $\frac{\sqrt{3}}{2}\hat{x}\pm\frac{1}{2}\hat{y}$ direction. In both cases, the distance between orbitals should be the same for a system without internal stress. However, in the superlattice of alternating graphene-fluorographene nanoribbons, the graphene channel is elongated along the longitudinal $\hat{x}$ direction, which increases the distance between orbitals that were initially in the $\frac{\sqrt{3}}{2}\hat{x}\pm\frac{1}{2}\hat{y}$ direction (odd $j$) if we compare it with the separation of orbitals along the $\hat{y}$ direction (even $j$), leading to weaker hopping terms.
@@
Figure 24: Variation of first-nearest neighbor matrix elements $t_{(j+1,j)}$ as a function of orbital position $j$. (a) 6-$\alpha\alpha$ system and (b) 8-$\alpha\beta$ system. (c) $t_{(j+1,j)}$ as a function of the inter-orbital distance ($d_{(j+1,j)}$). Lines represent the fitted functional dependence [indicated in (c)], with data points corresponding to the different system sizes.
@@
To quantify this effect, Fig. 24(c) plots $t_{j+1,j}$ against the interorbital distance ($d_{j+1,j}$) obtained from relaxed DFT geometries. The exponential dependence is apparent from the figure [30]. We fit this dependence using $t(d)=t_{0}\exp{(\gamma_{l}(\frac{d}{a_{0}}-1)})$, obtaining different parameters for even and odd $j$, reflecting the stress-induced anisotropy: for $j$ odd (black symbols), $t_{0}=-2.88$ eV and $\gamma=-3.16$; for $j$ even (gray symbols), $t_{0}=-2.82$ eV and $\gamma=-2.17$. In both cases, $a_{0}=1.42$Å. Importantly, the value of $t_{j+1,j}$ near the edges deviates from the fitted trends of the central region. This is due to the local distortion induced by the interface.
@@
In summary, to model $n$-$\alpha\alpha$ systems, we use the following system-size independent scheme for the first-nearest-neighbor hopping terms:
@@
The corresponding parameters for $n$-$\alpha\beta$ systems are
@@
In this case, the hopping term $t_{j+1,j}$ in the bulk of the channel has a constant value for $j$ odd, ${}^{1}t_{c}=-2.753$ eV, and a linear dependence, ${}^{1}\mathrm{t}_{t}(j)={{}^{1}}\mathrm{t}_{\alpha c}+\frac{{}^{1}\mathrm{t}_{\beta c}-^{1}\mathrm{t}_{\alpha c}}{2n-8}(j-4)$, for even $j$, with ${}^{1}t_{\beta c}=-2.854$ eV and ${}^{1}t_{\alpha c}=-2.804$ eV, in accordance with Fig. 24(b).